# CAMeleon: Interactively Exploring Craft Workflows in CAD

### v1

@@ 6.1. Interface Design
CAMeleon runs in user’s browser to maximize support across systems and architectures. The core of the interface is the way users explore different fabrication workflows, this is fundamentally supported by the comparison tool and the ability to browse different workflows. The other core aspect of the interface are the interactive guides to create human-centred CAM workflows. We discuss each individually in this subsection.
@@ 6.1.1. Comparison Windows
The CAMeleon comparison tool is designed to compare different workflows. The comparison tool lets users drag up to four workflows into the view-port, drawing inspiration from visual studio code editors. All workflows load the same 3D model, enabling users to clearly see and compare the potential outcomes side by side. This visual comparison helps users better understand how different workflows will influence the final product, making it easier to select the most suitable option based on their needs. Each window can be individually closed using the (x) icon on the top-right to free up space for other workflows or use more screen estate for the chosen workflow.
@@ 6.1.2. Workflow Browser & Filtering
To be able to compare different workflows, CAMeleon has a workflow browser panel on the left of the view-port. This is where users select workflows to be used in their fabrication process. Figure 7 shows different ways of using the browser. The most straight-forward is to scroll through the processes, however this obviously does not scale as the number of processes increases. To help them search more effectively CAMeleon lets them filter by keywords or by practical process-related dimensions.
Before categorizing the workflows in our survey, we placed them on different dimensions to compare and organize. These dimensions(e.g., what materials they use, what types of structure, or the degree of load-bearing capacity) formed an additional filter mechanism to find similar human centered CAM workflows. We found three categories of dimensions which are presented to users as shown in Figure 7(b).
Product-related Dimensions: These dimensions encompass factors such as load-bearing capacity, high-temperature tolerance, the need for a lightweight structure, and the level of fidelity required to reproduce intricate details. The choice of different support structures or materials can introduce unique features to the final product, enabling users to tailor their projects to specific needs.
@@
Structure-related Dimensions: These dimensions focus on the use and construction of support structures, jigs, or molds. They address aspects like the removability of the support structure, its integration within the object, flexibility, modularity, and reusability. This categorization helps users understand and manage the specific requirements and characteristics of the support structures used in their fabrication process.
@@
Machine-related Dimensions: The choice of fabrication machinery is closely linked to the selection of manufacturing processes. These dimensions guide users in identifying and selecting workflows based on the machines required, such as 3D printers, laser cutters, or wire benders.
6.1.3. Interactive Guide for human-centred CAM workflows
@@
As mentioned before, traditional CAM tools focus on machine-instructions, and fail to deliver instructions to humans who use the resulting artifacts and/or control the machines. We therefore propose human-centred CAM workflows which contains machine-instructions complemented with interactive guidance for users.
@@
Figure 8 shows such guidance, click through the different steps which allows them to follow the assembly process at their own pace and go back if they missed a step. (a) This assembly guide includes textual information, detailing how to assemble each component, with tracks and tips highlighted to ensure clarity. (b) Additionally, the guide includes links to external resources, such as YouTube tutorials, for in-depth guidance on the specific technique or examples of similar processes. This interface is designed to be similar to Instructables pages to slot in with existing experience of craftspeople and makers.
6.2. Implementation
CAMeleon is an open-source tool using Node.js and Three.js for 3D visualization and mesh processing. Three.js is used for its efficient handling of 3D models in the browser, while Node.js ensures scalable back-end support for file management and workflow coordination.
The tool processes STL files (containing triangle meshes), modifying the mesh and generating outputs like STL, SVG, or CSV depending on what machine is used. The system follows a modular architecture, where each workflow is independently defined in its own HTML file, allowing new workflows to be added without altering the core structure. The index.js file acts as a controller, dynamically loading and executing workflows based on user input. Each workflow follows a standardized pattern: receiving an STL file, processing it through Three.js libraries, and exporting the desired format.
6.3. Use Cases
We designed CAMeleon to work for three different categories of usage: during the design process, after the design process, and after fabrication.
During Design. During an ongoing design process, in CAD tools like Fusion360, users may want to explore how their design decisions impact the final fabrication workflow, or how different workflows can guide them in their design. To support this, we created a Fusion360 plugin as shown in Figure 9 that communicates with CAMeleon and displays previews of the resulting fabrication workflows. The browser-based nature of CAMeleon makes it flexible to be integrated with modern CAD tools.
After Design. CAMeleon takes volumetric models as input, these can come from a variety of sources, one of course is a CAD model when a user is done with their basic design. CAMeleon lets them design the model agnostic to the specific fabrication workflow, just specifying its volume. They can then use CAMeleon to generate and explore different human-centered CAM workflows. They could similarly download models shared online, or use generative AI tools (such as GENIE, Meshy, TRIPO), which are specifically good at generating workflow-agnostic shape. In both cases CAMeleon allows the generic volume to be converted to practical human-centered CAM workflows.
After Fabrication. Finally, CAMeleon can serve as a bit of a re-compiler of existing models. Users may like the shape of an object they see around them, 3D scan it, and then import that geometry in CAMeleon. The result is that they can re-fabricate the same shape but chose a different fabrication process instead. Figure 10 shows an example workflow of re-fabricating a Jellcat fox.
7. extending CAMeleon
To validate the extensibility of CAMeleon, we took two distinct approaches: implement expert workflows, and capture workflows presented in research papers as human-centered CAM workflows in CAMeleon. We approached six expert craftspeople, with their own bespoke workflows, to collaborate with us in implementing their workflows as extensions of CAMeleon. After we finished the implementation, we followed-up by reflecting with them on the results. We also identified three research papers describing specific fabrication workflows and implemented those in CAMeleon. In this section we describe the resulting workflows and the process of extending CAMeleon to encapsulate these workflows.
7.1. Collaborating Craftspeople
We collaborated with six craftspeople, (Table 1 shows their demographics and background), with the aim to re-implement their workflows in CAMeleon. Beforehand we conducted informal interviews with them after showing our software to better understand their perspective and workflows.
@@ 7.1. Collaborating Craftspeople
P2 | M | 4 years | Metal electroplating craftsman | Wire copper electroplating |
@@
P5 | F | 7 years | Papier-mâché craftsman | Paper mache shaping |
@@
7.2. Informal Interviews with Craftspeople
We draw the following observations from the informal semi-structured interviews.
@@ 7.2. Informal Interviews with Craftspeople
Four of the six craftspeople serve as educators in maker spaces. Their insights highlighted the value of integrating thinking around tool exposure and material understanding into the platform regarding workflow selection. P4 emphasized the importance of helping users understand the variety of tools available: ”The tools that people use are crucial. I think you’ve represented some, like laser cutting, mold making, needle felting, and styrofoam cutting… many students have no idea what tools they can use to shape or fabricate.” P3 appreciated how CAMeleonbalances tool affordances with material constraints: ”First and foremost, it’s about understanding the affordances and constraints of the tool and material we’re using… For example, if you’re using an Ultimaker, you’re limited by the build size… If a student or someone new to 3D printing starts from there, they won’t expect to print something 8 feet long.”
@@ 7.3. Implementing Craft Workflows
Here we lay out the process of re-implementing the workflows of our collaborating craftspeople in CAMeleon. We specifically discuss what aspects of the foundational workflows we managed to re-use in these workflows, highlighting valuable reusable modules. Figure 11 shows the reuse of modules of the original foundational workflows in creating the six additional craft workflows.
@@
While the re-use of existing modules made up the bulk of the code, the largest amount of time went into creating good instructions for the human-centered CAM workflows. This can typically not be reused from existing practice. Our collaborators had most of the required documentation readily available because some of them share their workflows online on platforms like Instructables. But materials existed in various levels of quality, and sometimes they tend to skip steps. We therefore re-recorded a series of interactive steps for each of the workflows. For practical extension, we added a feature in CAMeleon to link to a video or external page instead of reproducing all steps for each workflow. This reduces the barrier to entry for people with bespoke workflows.
@@
Figure 12 presents the re-implemented craft workflows in CAMeleon, to demonstrate that these work, (b) we exported and (c) fabricated a Stanford Bunny with each of these workflows. From left to right, the processes are Wire copper electroplating, Clay Sculpting, Silicone Molding, Needle Felting, Epoxy Coating, and Paper Mache Shaping.
@@
After implementation, we validated these workflows by showing them to the craftspeople and tuning basic parameters in the design based on their experience. The final workflows will be released as standard modules in CAMeleon with the craftspeople credited for their input.
7.4. Re-implementation of Research Workflows
To demonstrate extensibility of CAMeleon beyond classical craft workflows, we re-implemented three workflows we found in research papers. Again each of them reusing modules from our foundational workflows. With our template, we were able to quickly adapt workflows of Escape Loom (Deshpande et al., 2021), StackMold (Valkeneers et al., 2019), and Bamboo Agents (Gao et al., 2023) by adjusting parameters in key modules such as the slice generator, process convex hull, mesh intersection line, and export tools (STL, SVG). The flexibility and modularity of our system allowed for efficient adaptation with minimal changes, showcasing how the re-usability of foundational modules streamlines the integration of diverse digital fabrication workflows.
However, during the implementation process, we discovered that while the software behind these workflows excels at handling geometric calculations, it falls short in accurately recording the full craft process. Although some steps have been simplified through digital tools, the workflows are still incomplete in capturing the nuances of hands-on craftsmanship. Key aspects of fabrication, which also rely on experiential knowledge, remain outside the scope of what these tools can document. While they help reduce certain challenges, they cannot fully represent the entire craft workflow, leaving gaps in the process that only skilled craftsmen can bridge.
8. USER STUDY
To evaluate the usability of an end-to-end experience with CAMeleon, we conduct a user study.
Participants We recruited 12 participants from our university (number of female students = 8, age mean = 24.66 years, SD = 1.66). Most participants (9) reported having beginner proficiency with CAD tools.
Task. Participants were tasked to fabricate a physical model from a digital design created with Luma AI (a generative AI tool). After importing their AI-generated 3D model into CAMeleon, they explored different ways of fabricating their object and used the respective human-centered CAM workflow to fabricate their object.
Metrics. We measured to general usability experience of using CAMeleon with questions from the standard SUS survey (Brooke, 1996). This survey was presented to participants after they exported their model from CAMeleon. We furthermore conducted semi-structured interviews a day later after participants fabricated their final object.
Procedure. A member of the research team started each session by outlining the research objectives and introducing CAMeleon. This introduction took approximately 5 minutes. Participants were given 10-minute to outline what they intend to fabricate. Participants generated their target model and used CAMeleon to explore and determine their target workflow. When done, they answer a short questionnaire aimed at capturing their initial reactions and usability experiences with CAMeleon. Researchers used the exported machine-instructions to fabricate the parts over night, the following day, participants returned to manually assemble the parts using their chosen human-centered CAM workflow with interactive guidance from CAMeleon.We concluded with a 15-minute semi-structured interview to gather more detailed feedback about their interactions with CAMeleon and its effectiveness in assisting them from digital model to physical assembly.

### v2

@@ 6.1. Interface Design
CAMeleon is a browser tool. Users compare workflows using the comparison tool and the ability to browse different workflows. Once a workflow is selected, the other core aspect of the interface is the interactive guides to create human-centered CAM workflows. We discuss each individually in this subsection.
@@ 6.1.1. Comparison Windows
The comparison tool (shown in Figure 6) lets users drag up to four workflows into the view-port. All workflows load the same 3D model, enabling users to compare potential outcomes side by side. In addition to visual comparison, the comparison windows also include information about the processes involved in each workflow.
Users can preview key manual steps and required machines, as outlined in the human-centered CAM workflows. This feature allows workflows to be compared not only by their visual results but also by the steps and resources they require. Users can close workflows to compare with others or free up space to focus on 1-2 workflows in more detail.
@@ 6.1.2. Workflow Browser & Filtering
CAMeleon has a workflow browser panel on the left of the interface (demonstrated in Figure 7). The most straight-forward is to scroll through the workflows, however this does not scale as the number of workflows increases. Users therefore can filter by keywords related to the specific workflow or meta-properties such as workflow criteria, machines, or materials.
Using the workflows from our survey, we identified three categories of workflow-related dimensions which are presented to users as shown in Figure 7(b).
Product-related Dimensions: Product properties such as high-temperature tolerance and the fidelity required to reproduce intricate details.
@@
Structure-related Dimensions: The use and construction of support structures, jigs, or molds. This categorization helps users understand and manage the specific requirements and characteristics of the support structures used in their fabrication process.
@@
Machine-related Dimensions: The choice of fabrication machinery is closely linked to the selection of fabrication workflows. These dimensions guide users in identifying and selecting workflows based on the machines required, such as 3D printers, laser cutters, or wire benders.
6.1.3. Interactive Guide for human-centered CAM workflows
@@
As mentioned before, traditional CAM tools focus on machine-instructions, and fail to deliver instructions to humans who use the resulting artifacts and/or control the machines. We therefore propose human-centered CAM workflows which contains machine-instructions complemented with interactive guidance for users.
@@
Figure 8 shows such guidance, users click through the steps one at a time. (a) This guide includes textual information, detailing how to assemble each component, with tips highlighted to ensure clarity. (b) Additionally, the guide includes links to external resources, such as YouTube tutorials, for in-depth guidance on the specific technique or examples of similar processes. These guides are currently not specific to the design of the user, they serve as a general explanation of the workflow, future work could explore how to leverage dynamic rendering techniques to contextualize these instructions for the model at hand.
6.1.4. Loading tool
The loading tool enables the import of STL models. However, when users adopt a model for a specific fabrication workflow outside CAMeleon during the design phase, the model may be constrained to a specific workflow (say, a structure of laser-cut plates). CAMeleon’s import recovery aims to reconstruct a volumetric representation, enabling users to explore alternative workflows again. Two reconstruction methods are supported: Voxelization with Poisson surface reconstruction and Alpha shape reconstruction, each addressing different structural characteristics.
Figure 9 illustrates the application of these two methods using the Stanford Bunny model. For a laser-cut interlocking planar structure (Figure 9a), Alpha shape reconstruction accurately recovered the inner geometry. For this wire-bending structure (Figure9 b), Voxelization with Poisson surface reconstruction produced a smooth and complete recovery of the outer surface. However, if a model is already volumetric, there is no need for pre-processing other than basic mesh repair (filling holes, completing triangles). The preview of the different imports gives users a sense of what their model will look like in CAMeleon as a result of the different pre-processing algorithms.
7. Implementation
CAMeleon is built on Node.js, Three.js (using WebGL) and Python. Three.js serves as the foundation, providing 3D rendering and geometry handling in the browser. Building on this, CAMeleon extends Three.js with CAD-specific geometry processing, contour simplification, projection generation, and visualization tools (e.g., draft angle analysis, bounding box helpers, and color-coded feedback). It supports output formats like STL, SVG, and CSV, tailored to different machines. The Python-based back-end complements these features by handling computationally intensive geometry operations, ensuring efficient processing.
CAMeleon’s workflows are implemented using a modular architecture. Each workflow leverages reusable modules to address tasks, such as mold making, foam cutting, or interlocking structure generation. By remixing these modules, developers compose new workflows as demonstrated in the ”Extending CAMeleon” section.
As shown in the Figure 10, When users load a new model, the WorkflowController calls the Model Loading for the specialized task of 3D model loading and pre-processing. Once the model is loaded, the front-end communicates with the Flask-based back-end through REST APIs. These APIs facilitate file upload and initiate geometry operations, leveraging libraries like Open3D and PyMeshLab for computationally intensive tasks such as model simplification, Alpha Shape generation, and voxelization.
7.0.1. Workflow Modules
CAMeleon modules are organized into four categories and listed in Table 1, covering the complete workflow for 3D visualization and fabrication:
Model Loading. Handles the import and initial processing of 3D models, including STL file loading and basic geometry adjustments.
Geometry Processing. Performs core computational tasks on 3D geometry, such as intersections, contour simplifications, and projections for specific workflows.
Visualization. Enhances understanding of 3D models through tools like draft angle analysis, dimension labeling, model information helpers, and color-coded feedback for manufacturing validation.
Export. Converts processed models or data into formats such as STL, SVG, or CSV for fabrication workflows.
7.0.2. Distribution
CAMeleon is available as a simple browser tool on RELEASED ON ACCEPTANCE, and the source code can be accessed via RELEASED ON ACCEPTANCE.
Category | Module | Description |
Model Loading | Model Loading | Load and process STL files, with options for centering, scaling, mesh rotation (x/y/z), and bounding box handling. |
Geometry Processing | Hollow Triangle Gen | Generate hollow triangles from vertices, faces, and normals. |
Threaded Rods Gen | Create 2D and 3D shapes for threaded rods. | |
Tube Gen | Generate tubes with specified color, opacity, and radius. | |
Interlock Joints Gen | Create interlocking joints based on points and dimensions. | |
Split Volume | Create positive and negative molds with air vents. | |
Intersect | Generate intersections and contours using meshes and planes. | |
Extrude | Generate 3D extruded shapes from lines. | |
Extrude to Direction | Extrude shapes in a specific direction with defined depth. | |
Loft | Create lofted surfaces from inner and outer contours. | |
Calc Draft Angles | Analyze draft angles with color mapping. | |
Project Mesh | Project meshes into 2D lines. | |
Simplify Geom | Simplify geometry based on distance and angle thresholds. | |
Calc Centroid | Compute the centroid of triangular vertices as a Vector3. | |
Calc Line Inters | Calculate intersection points for lines in specific planes. | |
Vertex Scaling | Scale vertices around a centroid using a scale factor. | |
Visualization | Model Info Vis | Display mesh details and properties. |
Visualize Mold | Visualize positive and negative mold parts. | |
Tutorial & Info | Provide text, images, and links for user guidance. | |
Color Legend | Show visual legend for angle ranges and other properties. | |
3D Visualization | Render meshes in 3D. | |
Geom Visualization | Visualize points and lines with legends. | |
Export | Export SVG | Save lines as SVG files. |
Export CSV | Save lines as CSV files. | |
Export STL | Save meshes as STL files. |
8. Expert Evaluation of CAMeleon with craftspeople
To validate CAMeleon’s ecological validity, we interviewed and collaborated with six expert craftspeople to replicate their bespoke workflows inCAMeleon, we reflect with them on the resulting workflows. This section outlines the resulting workflows and the process of extending CAMeleon to support these workflows.
8.1. participants
Table 2 shows the demographics and background of our collaborating expert craftspeople and the workflows they choose to integrate into CAMeleon.
@@ 8. Expert Evaluation of CAMeleon with craftspeople
P2 | M | 4 years | Electroplating | Wire copper electroplating |
@@
P5 | F | 7 years | Papier-mâché craftsman | Papier-mâché shaping |
@@
8.2. procedure
We approached this in three phases: first we showed participants CAMeleon and conducted semi-structured interviews to understand their perspective and workflows. We then proceeded by implementing their workflows in CAMeleon and follow-up with an interview to reflect and collect their feedback.
8.3. Initial Interviews
We draw the following observations from the initial interviews.
@@ 8.3. Initial Interviews
Four experts also serve as educators. Their insights highlighted the value of integrating thinking around tool exposure and material understanding into the platform regarding workflow selection. P4 emphasized the importance of helping users understand the variety of tools available: ”The tools that people use are crucial. I think you’ve represented some, like laser cutting, mold making, needle felting, and styrofoam cutting… many students have no idea what tools they can use to shape or fabricate.” P3 appreciated how CAMeleon balances tool affordances with material constraints: ”First and foremost, it’s about understanding the affordances and constraints of the tool and material we’re using… For example, if you’re using an Ultimaker, you’re limited by the build size… If a student or someone new to 3D printing starts from there, they won’t expect to print something 8 feet long.”
@@
8.4. Insights to support workflows in CAMeleon
The interviews provided valuable insights into how craftspeople consider workflows, support entry-level makers, and balance traditional crafting with digital fabrication.
A frequent theme was balancing dimensions like tool affordances, surface quality, and material constraints when selecting and comparing a workflow. Participants emphasized that providing clear and accessible guidance not only empowers users to navigate these considerations effectively but also instills in them a sense of confidence and capability to make informed choices about their projects.
Another key takeaway was the immense value of experiential learning for novices. Rather than simply comparing workflows, participants highlighted the need to provide users with essential information to deepen their understanding of tools and materials. This approach, which emphasizes hands-on experience, lets users quickly try different methods, accelerating their learning and supporting hands-on exploration. These discussions underscore the importance of designing learning platforms that not only convey technical skills but also cultivate an experimental mindset, encouraging users to appreciate both the ease of digital fabrication and the tactile, sometimes unpredictable, nature of traditional craft methods.
The interviews also revealed the dynamic interplay between traditional craftsmanship and digital fabrication. While digital tools such as CAD provide efficiency and precision, the tactile understanding developed through hands-on craftsmanship remains crucial. Participants emphasized how integrating these two realms opens new creative possibilities, allowing users to transition between designing in the digital space and physically crafting their projects.
@@ 8.5. Implementing Craft Workflows
We lay out the process of implementing the workflows of our collaborating craftspeople in CAMeleon. This required close collaboration with the craftspeople to understand their bespoke practices and translate them into CAMeleon worfklows.
@@
The most time-consuming aspect was creating good instructions for human-centered CAM workflows. This can typically not be reused from existing practice. Our collaborators had most of the required documentation readily available. But documentation existed in various levels of quality, and sometimes they skip steps. We therefore recorded a series of steps for each of the workflows. For practical extension, we added a feature in CAMeleon to link to a video or external page instead of reproducing all steps for each workflow. This reduces the barrier to entry for people with bespoke workflows. A complete example of a human-centered CAM workflow for Needle Felting is provided in the appendix to illustrate the detailed steps and instructions.
@@
Figure 11 presents the re-implemented craft workflows in CAMeleon, to demonstrate that these work, we exported and fabricated a Stanford Bunny with each of these workflows. From left to right, the processes are Wire copper electroplating, Clay Sculpting, Needle Felting, Papier-mâché Shaping Silicone Molding, and Epoxy Coating.
@@
After implementation, we validated these workflows by showing them to the craftspeople. The workflows will be standard modules in CAMeleon with the craftspeople credited for their input.
8.6. Feedback
After seeing their workflows in CAMeleon, craftspeople highlighted the importance of including detailed guidance in Human-Centered CAM workflows, which they saw as crucial for supporting users. P5 emphasized the importance of explaining material choices and techniques, sharing insights on working with pulp and adhesives: “If you add too much glue and grind the paper too finely, it becomes unmanageable… Once it dries, you can mix the pulp and glue like clay to shape it.” Similarly, P6 offered insights on silicone mold workflows, discussing disassembly methods and material selection, such as, “Two-part molds reduce breakage risks… whereas single molds might crack more easily at sharp corners.” P4 suggested adding more detailed guidance on tool selection, such as ”If you needed to cut something, what type of saw would you use, or would you use a chisel? Or would you use a drill or jigsaw?…It has real teaching potential, showing users what’s possible and providing an estimate of how the process will flow.” Human-Centered CAM was praised for its ability to bridge computational design, digital fabrication, and physical workspaces. P1 noted that CAMeleon encourages users to think about physical structures and manufacturing during the design phase: “It allows users to integrate modeling with considerations for physical structures and fabrication.”
CAMeleon’s educational potential stood out to participants. P4: “It’s great for first-time users or people without sculptural experience.” P4 appreciated the diversity of techniques showcased—such as foam cutting or papier-mâché—stating, “You’re presenting options users might never have considered.” P3 emphasized how CAMeleon fills gaps in curricula:“The curriculum is often missing basic design and making knowledge. Your software helps guide novice makers to recognize the value of exploration.” Furthermore, P3 added: “We have advanced tools, but we’re hesitant about novice makers using them. Your software fills that gap, helping students gain both access and understanding.”
Analysis feature like the draft angle analysis received acclaim. P6: “It identifies weak areas—if there’s too much blue, the mold design might need adjustments.” P6 also suggested enhancing computational capabilities, such as adding gates to support delicate mold areas: “For example, supporting the ears of a chocolate rabbit mold can reduce breakage risks.”
9. Technical evaluation on the Extensibility of CAMeleon
To do a deeper dive into the extensibility of CAMeleon, we implemented workflows described in three research papers. Traditionally, research workflows focus on their core academic contributions, and thus can be hard to replicate. We demonstrate how the modular architecture of CAMeleon supports this process and thus creates a platform for experimentation across workflows beyond bespoke crafts.
9.1. Selected workflows
EscapeLoom (Deshpande et al., 2021), StackMold (Valkeneers et al., 2019), and Bamboo Agents (Gao et al., 2023) were selected as representative workflows due to their alignment with the overarching theme of integrating traditional crafts with digital fabrication tools. Each of these works, sourced from the ACM Digital Library under the Human-centered computing Index Terms, emphasizes interactive systems and tools that facilitate novel approaches to craft-making. The decision to replicate these workflows stems from their focus on enhancing user creativity and expanding the possibilities of fabrication through the interplay of manual craftsmanship and computational methods.